Accelerator Update for February 4, 2009 - February 6, 2009 |
Wednesday Feb 4 The day shift began with Operations monitoring the Tevatron (TeV) store 6771 (71.32E30) and the Recycler stash (315.0E10), with the Antiproton Source stacking (31.2E10), with the Linac, Booster, and Main Injector (MI) operating normally, with the Collider Detector at Fermilab (CDF) and D-Zero (D0) experiments using the colliding beams, and with the MiniBooNE and NuMI experiments taking beam. Operations terminated store 6771 at 10:44 AM. Operations put the TeV into shot setup at 10:55 AM. Operations established store 6772 at 12:18 PM with an initial luminosity of 321.07E30. Linac and Booster were turned off at 1:00 PM for RF repairs. Stacking and beam to MiniBooNE resumed at 2:25 PM. Beam returned to NuMI at 3:21 PM. The Tevatron quenched at 4:00 PM, Operations and experts investigated. Accesses into the Tev began at 6:04 PM to repair the cryostat vacuum at A3. CDF and D0 used the opportunity to access their collision halls. CDF came out of access at 7:16 PM D0 came out of access at 8:27 PM. Vacuum pump down at A3 began at 9:13 PM. At 11:25 PM, vacuum at A3 is getting worse, Cryo and Mechanical Techs are accessing the Tev. Thursday Feb 5 The midnight shift began with Operations monitoring the stash (257.9E10), with Pbar stacking (25.5E10), and with NuMI and MiniBooNE taking beam, and with the Tevatron in access. Tevatron access was completed at 2:37 AM. Tevatron turn on and checkout began. D0 and CDF went into access at 7:00 AM. CDF and D0 accesses were completed at 9:10 AM. The Tev completed a wet squeeze at 12:37 PM, orbit corrections were made at this time. Operations put the TeV into shot setup at 1:34 PM. Operations established store 6778 at 3:44 PM with an initial luminosity of 272.26E30. Booster RF station #18 (BRF18) failed at 6:35 PM. BRF18 was bypassed and beam returned at 6:48 PM. Friday Feb 6 The midnight shift began with Operations monitoring store 6778 (94.6E30) and the stash (317.7E10), with Pbar stacking (23.0E10), and with NuMI and MiniBooNE taking beam. Operations terminated store 6778 at 4:10 AM. Operations put the TeV into shot setup at 4:23 AM. The Tevatron quenched at 5:57 AM, Operations and experts investigated. Beam was turned off at 7:30 for Booster RF work, Pbar will also use this access time. Tev access began at 8:00 AM, CDF and D0 will also access. The Plan for today and the Weekend Recover from the Tevatron quench, complete the Booster and Pbar accesses and turn machines on. Return to stacking, stashing and storing. |
|
More Information Comments and Suggestions |
| last modified 02/06/2009 email Fermilab |
| Security, Privacy, Legal | |